Multinational Financial Management Exercises Question 1. Spot rate: Gi ven the following direct quotes, calculate the equivalent indirect quotes. A. $0.09/Mexican pesos B. £0.75/E C. Rupees 30.50/ CS Question 2. spot rate: Suppose a BMW 745i car is priced at S60,000 in New York and 50,000 in Berlin. In which place is the car more expensive if the spot rate is $1.23/? Forward rate: GWC Corporation has contracted with an Indian technology firm for hardware products. The payment of 10,000,000 rupees (Rs.) is due in 60 days. How much will the cost be in dollars if the 60-day Question 3. forward rate is Rs. 50.25/S? Question 4. Cross rate: Barney Bank has offered the following exchange rate quotes: *214/E and Korean won 15%. What is the cross rate between the Korean won and the British pound? Question 5. Cross rate: Jensen Corporation observes the Swiss franc (SF) is currently quoted at €0.825 F and the Sweden krona (SK) ?s quoted at €0.125/SK. What is the SK/SF cross rate? Multinational Finance Spreadsheet Example ASSIGNMENT # NAME: LAB TIME: Answer Sheet for Multinational Finance Exercises: Question 1A Qestion 1B Question 1C Question 2 Dollar cost in Berlin $ The cost of the car is higher in Question 3 Question 4 Question 5
Answer #1

1A.

Indirect Quote = 1 / 0.09

= 11.11 Mexican Peso Per US dollar

Indirect quote is 11.11 Mexican Peso Per US dollar.

1B.

Indirect Quote = 1 / 0.75

= 1.33 Euro Per Pound

Indirect quote is 1.33 Euro Per Pound.

1C.

Indirect Quote = 1 / 30.75

= 0.0325 Rupee per Canadian Dollar

Indirect quote is 0.0325 Rupee per Canadian Dollar.

2.

Price of Car in US = $60,000

Price of car in Germany = 50,000 Euro

Exchange rate = $1.23 per euro.

Price of Car in Germany in dollar = 50,000 × 1.23

= $61,500

Price of Car in Germany in dollar is $61,500.

So cost of Car in Germany is higher than US.